# Fisher information flow in artificial neural networks

### v1

@@ VII Acknowledgments
We thank Sabine Andergassen and Thomas Gärtner for helpful discussions. L.M.R. was supported by the Austrian Science Fund (FWF) through Project No. P32300 (WAVELAND). I.S. and D.F. acknowledge financial support from the UK Engineering and Physical Sciences Research Council (EPSRC grants No. EP/T00097X/1, EP/Y029097/1). D.F. acknowledges support from the UK Royal Academy of Engineering Chairs in Emerging Technologies Scheme. The authors acknowledge the TU Wien Bibliothek for financial support through its Open Access Funding Programme.
@@ A.1 Sufficient statistics and conservation of Fisher information
Sufficient statistics [23, 10, 45] are transformations that preserve all the relevant features of the data with regard to the parameter of interest. Without much loss of generality we state that sufficiency of a statistic and preservation of FI are equivalent (in [78] a counter example is constructed, which is, however, not relevant here). Finding a sufficient statistic is itself not a difficult task, as any bijective transformation preserves information. In parameter estimation, however, an additional desired property is to simultaneously compress the data to a dimension that coincides with the number of parameters to be estimated. Unless the data follows an exponential family of distributions, a non-trivial sufficient statistic that also compresses the data does generally not exist, hence one cannot expect that it is possible to compress the data without losing FI.[22] Motivated by [10, 62, 63, 59] we relax the conditions of sufficiency and we speak of conservation when the FI stays approximately the same as the data is transformed. Whether it is possible to approximately conserve all of the FI depends on the estimation task as discussed in the following.
@@
FI can be lost whenever the function connecting two layers of a ANN cannot be inverted. This happens if either the dimension of the data is reduced or if the activation function [40], i.e. the non-linearity that acts upon all hidden layers’ neurons, is not bijective. Training a neural network to estimate a single parameter from image data requires that the data is projected from $d$ to one dimension, where the former represents the number of pixels in the image. Hence, the network cannot always represent a sufficient statistics and one cannot expect that all of the FI present in the data is conserved as the data is processed by the neural network. Whether a model can reach the (bias-corrected) CRLB and conserve all of the FI depends on the distribution and the signal-to-noise ratio [43, 79] but also on the range of parameters that the model is trained on [80, 58]. The broader this range is the harder it becomes for the model to reach the CRLB and to conserve the FI. One way to understand this behavior is to note that the CRLB is determined by the log-likelihood function and it’s first derivative with respect to the parameter. If the log-likelihood function has a complicated shape and one represents it as a Taylor series on an extended parameter interval, higher order derivatives need to be taken into account and the CRLB can be too lose [80]. If the parameter range on which the neural network is trained is sufficiently small, however, these higher order derivatives contribute only little and the CRLB becomes (approximately) reachable.
@@
In order to investigate if the effects mentioned above play a role in the mismatch between input and output FI we observe for the experimental data, we train ANNs to estimate the position of the object, each ANN with a different number of positions shown during the training. The FI is always evaluated using only the $3$ central positions. The fraction of the input FI that arrives at the output of the ANNs trained on different parameter ranges is shown in Fig. A1.
@@
Figure A1: The fraction of the input FI that arrives at the output layer of an ANN versus the number of positions used to train the ANN. The spacing between the position is equal ($\Delta\theta=0.3$ µm), hence more positions correspond to a broader spatial range. $13$ positions corresponds to the parameter range we used to generate the results in the manuscript. We observe that the narrower the range, the more FI is conserved. The ANNs layers from input to output consist of $[432,1000,500,250,50,1]$ nodes.
@@
We clearly observe that for narrower parameter ranges, more FI is conserved and when only the central $3$ positions remain in the parameter range (those where we evaluate the FI), the mismatch disappears entirely. This result provides strong evidence that the mismatch we observed is due to the finite range of parameters we were working with.
@@
Note that there can be other reasons for such a mismatch. Examples are a (stochastic or deterministic) error in the estimated FI, a bias picked up by the estimator that cannot by taken into account with the parameter-resolution at hand or that not enough training data is available to achieve the best performance possible.
@@ A.2 Statistical properties of the estimated Fisher information
Since in our algorithm of LFI maximization we estimate the FI using a finite-sized data set, the estimated FI itself is a random variable. Here, we provide an analytical expression for the variance of this random variable, allowing to gauge whether the data at hand allows for a sufficiently precise estimation of the FI. Note that our estimate of the FI is the LFI of the data after it is projected into a higher dimension (see II.5). The variance of our estimate is thus the variance of this LFI of the transformed data. In order to approximate the variance of the LFI, we carry out a derivation similar to that in [53] and find the expression
@@
where $\sigma_{\hat{J}}^{2}$ is the variance of the estimated LFI $\hat{J}$, $J$ the LFI, $d$ the dimension of the random variable and we introduce $L=JN_{\mathrm{sample}}\Delta\theta^{2}$ with the sample size $N_{\mathrm{sample}}$. To estimate the FI we gradually increase $d$ until convergence is achieved. The results from the algorithm are precise if the fluctuations of the estimated LFI are small compared to the LFI, i.e., $\eta(L)\ll 1$ for the largest value of $d$ we chose in the algorithm. Note that if the LFI curve as a function of the dimension is not sufficiently smooth, we average over a few dimensions to smoothen the curve and identify the plateau more easily; it is reasonable to assume that this summation of LFIs does not increase the standard deviation.
From Eq. 8 we learn that the quantity $L$ should be large in order to precisely estimate the FI. This means that the sample size must be large enough to compensate for the squared finite-difference step size, which must be chosen small enough such that the discretization error remains small.
We also note that the term in Eq. 8 that is proportional to $d$ decreases with $1/L^{2}$ while the term independent of $d$ only decreases with $1/L$. Hence, for large $L$ the uncertainty of the estimated FI depends only weakly on the dimension of the projection.
@@ B.1 Optical setup
Broadband light is generated by a fiber optic illuminator (Oriel, Model 77501) coupled to a multimode fiber. Light is outcoupled and attenuated using a neural density filter (optical density 4) before being used to illuminate the sample. On this sample, a drawing of a “space invader” was engraved onto a chromium layer (CAD/Art Services). The position of the sample is controlled by a motorized stage (PI M-122.2DD1). Transmitted light is collected by a microscope objective (Olympus Plan N $\times$20, 0.4 NA) and imaged using a tube lens ($f=150$ mm) onto a camera (PCO Panda 4.2 sCMOS).
@@ B.2 Acquisition procedure
Images are measured for $13$ different positions of the sample, which is linearly translated along the $x$ direction with a step size of $0.3$ µm. The range of positions covered by the object is thus comprised between x=1.8x=-1.8 µm and $x=+1.8$ µm. The image dimension is set to $32\times 64$ pixels, with a binning procedure such that each pixel corresponds to the sum of $4\times 4$ physical pixels of the camera. The exposure time is set to $10$ µs, which results in a very low signal-to-noise ratio. 300 000300\,000 images are measured for the three central positions (x=0.3x=-0.3 µm, $x=0$ µm, and $x=+0.3$ µm), and 50 00050\,000 images are measured for all other positions. Finally, a set of images is measured for different sample positions using a long exposure time ($100$ ms) for illustrative purposes.
@@ B.3 Noise of the experimental data
Figure A2: Panel (a) shows histograms for a generic marginal distribution of the experimental data (a pixel at the center of the image). The number of histogram bins is $70$. We use a total of 50 00050\ 000 data points. Panel (b) shows histograms for a Gaussian distribution with the same mean and variance as the experimental marginal distribution, with the same number of data points and number of bins.
@@ C.1 Generation of the Gaussian data set for observing the Fisher information flow
80% of the data set is used to train the ANN while the remainder is reserved for validation. A separate data set is prepared for evaluating the FI flow. To approximate the derivative in the expression for the FI by a symmetric finite difference scheme, we first generate data on an equally spaced grid, but with three grid points $\theta\in\{-0.03,0,0.03\}$. 150 000 data points are generated per grid point.
@@ C.2 Generation of the experimental data set for observing the Fisher information flow
As described in Subsec. B.2 the training data set consists of 50 00050\,000 images for each of the $13$ positions of the scatterer. Observing that information about the scatterer only arrives at a central region of the camera, we crop the images symmetrically around the center, such that the size of the images is reduced from $32\times 64$ to $18\times 24$. This procedure has a small effect on the information content of the data but reduces the memory required to store the data sets. Before it is fed to the network, the data is standardized by subtracting the intensity averaged over all images and dividing by the corresponding standard deviation. This invertible transformation does not alter the information content. 80% of the data set is used to train the ANN while the remainder is reserved for validation.
@@
C.3 Generation of the Gaussian data set for observing Fisher information based early stopping
The data set is equivalent to that of Appendix C.1 but with $101$ parameter grid points and with $7000$ data points for each grid point and ranging from $\theta=-1.5$ to $\theta=1.5$.
C.4 Generation of the experimental data set for observing Fisher information based early stopping
@@ Appendix D Neural network architectures and training parameters
Table AI: Neural network architectures and training parameters for the ANNs employed to observe the flow of FI and early stopping based on FI, both for the Gaussian training data and the experimental training data.
@@ Appendix E Calculation of Fisher information
The LFI itself already yields satisfying results in most of the ANN’s layers. We employ the LFI maximization algorithm to find corrections to the LFI. Especially in early stages of the training there is no guarantee that the LFI is a good approximation to the true FI, thus the LFI maximization algorithm yields more accurate and reliable results compared to only using the LFI. The linear transformations for embedding into a higher dimensional space are random matrices with independent and identically distributed entries following a normal distribution with 0 mean and a standard deviation of $\sigma=1$. We then apply the non-linearity
@@
with $\alpha=0.7$ to each component of the transformed random variable and eventually add noise following a normal distribution with 0 mean and a standard deviation of $\sigma_{\mathrm{noise}}$. In each iteration we increase the dimensionality by $\Delta d$. As the criterion for convergence, we use that the difference between the smallest and the largest value of the LFI within the previous $3$ iterations must be smaller than that of the first $3$ iterations by a factor of $\kappa$. If during the first $3$ iterations the increase of the LFI is smaller than a fraction of $\gamma$ of the LFI with $d^{\prime}=d$, we stop the maximization procedure and choose the LFI averaged over the first $3$ estimates as the estimate of the true FI. The convergence criteria could not be fulfilled for all layers and epochs of the networks trained on the different data sets, thus the LFI itself is in these cases used as a rough approximation of the true FI. If the optimized LFI is below the LFI of the data set we use the LFI as the estimate for the FI. This case occurs whenever the LFI is very close to the FI and the algorithm slightly underestimates the FI due to the added noise. In Table AII we list the parameters of the LFI maximization algorithm together with the exceptions, where the algorithm did not converge.
@@
Table AII: Parameters used for calculating the FI with the LFI maximisation algorithm. Data is extracted at each layer of the ANNs that are trained to observe the FI flow.
@@
Note that it is not necessary to apply the LFI based algorithm to the input layer of the ANN trained with the Gaussian training data set as the FI of this data set is known analytically (see Subsec. II.4).
@@
Fig. A3 and Fig. A4 show the LFI as a function of $d^{\prime}-d$ for an exemplary layer of the network trained on the Gaussian data set and one of the network trained on the experimental data set. Note that the exact shapes of these curves depend also on the magnitude of the noise that is added for regularization.
@@
Figure A3: LFI during the LFI maximization for Gaussian data. The data is extracted at the second hidden layer before the first epoch of training. The curve depicts the estimated LFI for consecutive embeddings. The dimension of the original data set is $d=100$ and it is increased in steps of 30.
@@
Figure A4: LFI during the LFI maximization for the experimental training data set. The data is extracted at the output layer as the network is fully trained. The curve depicts the estimated LFI for consecutive embeddings. The dimension of the original data set is $d=1$ and it is increased in steps of 30.
Appendix F Parameter resolved flow of Fisher information
@@
As the FI is local in the parameter, i.e., it can be different for different values of $\theta$, we can access the local flow of information through an ANN. To demonstrate this, we train an ANN on the Gaussian data set where the estimated parameter is proportional to the mean and calculate the layer-wise FI as a function of this parameter. The neural network architecture and the specifics of the training data set are the same as those we used to observe the FI flow. We evaluate the FI at $15$ points in the range $\theta\in\{-1,1\}$ and generate for each point $3$ samples of size 100 000100\,000, one at the point of observation $\theta$, one at $\theta-0.1$ and one at $\theta+0.1$. The result is shown in Fig. A5.
@@
Figure A5: Fisher information flow through an ANN trained on the Gaussian data set. The curves show the (normalized) Fisher information as a function of the estimated parameter $\theta$. Each curve corresponds to a certain layer of the ANN, which is indicated by the color of the curve. The network architecture is the same as that for the flow of FI for a single parameter. Note that we stop the training at an early epoch such that it is possible to distinguish the different curves. The FI is thus not conserved in this example.
@@
We observe that the flow of FI is a highly non-uniform and non-symmetric function of the parameter. The ANN thus learns to process data at different parameters differently. The parameters we use for our algorithm for calculating the FI are the same as in Table AII, except for $\sigma_{\mathrm{noise}}=0.05$.
@@
Appendix G Testing our methods on data following a correlated Lognormal distribution

### v2

@@ VII Acknowledgments
We thank Sabine Andergassen and Thomas Gärtner for helpful discussions. L.M.R. was supported by the Austrian Science Fund (FWF) through Project No. P32300 (WAVELAND). I.S. and D.F. acknowledge financial support from the United Kingdom Engineering and Physical Sciences Research Council (EPSRC Grants No. EP/T00097X/1 and No. EP/Y029097/1). D.F. acknowledges support from the United Kingdom Royal Academy of Engineering Chairs in Emerging Technologies Scheme. The authors acknowledge the TU Wien Bibliothek for financial support through its Open Access Funding Programme.
VIII Data availability
The experimental data [78] and an exemplary code for processing the experimental data and the numerically simulated data [79] are publicly available.
@@ A.1 Sufficient statistics and conservation of Fisher information
Sufficient statistics [23, 10, 45] are transformations that preserve all the relevant features of the data with regard to the parameter of interest. Without much loss of generality we state that sufficiency of a statistic and preservation of FI are equivalent (in Ref. [80], a counterexample is constructed, which is, however, not relevant here). Finding a sufficient statistic is itself not a difficult task, as any bijective transformation preserves information. In parameter estimation, however, an additional desired property is to simultaneously compress the data to a dimension that coincides with the number of parameters to be estimated. Unless the data follow an exponential family of distributions, a nontrivial sufficient statistic that also compresses the data does generally not exist; hence, one cannot expect that it is possible to compress the data without losing FI [22]. Motivated by Refs. [10, 62, 63, 59], we relax the conditions of sufficiency and we speak of conservation when the FI stays approximately the same as the data are transformed. Whether it is possible to approximately conserve all of the FI depends on the estimation task as discussed in the following.
@@
FI can be lost whenever the function connecting two layers of an ANN cannot be inverted. This happens if either the dimension of the data is reduced or if the activation function [40], i.e., the nonlinearity that acts upon all hidden layers’ neurons, is not bijective. Training a neural network to estimate a single parameter from image data requires that the data are projected from $d$ to one dimension, where the former represents the number of pixels in the image. Hence, the network cannot always represent a sufficient statistics and one cannot expect that all of the FI present in the data is conserved as the data are processed by the neural network. Whether a model can reach the (bias-corrected) CRLB and conserve all of the FI depends on the distribution and the signal-to-noise ratio [43, 81] but also on the range of parameters that the model is trained on [82, 58]. The broader this range is the harder it becomes for the model to reach the CRLB and to conserve the FI. One way to understand this behavior is to note that the CRLB is determined by the log-likelihood function and its first derivative with respect to the parameter. If the log-likelihood function has a complicated shape and one represents it as a Taylor series on an extended parameter interval, higher-order derivatives need to be taken into account and the CRLB can be too lose [82]. If the parameter range on which the neural network is trained is sufficiently small, however, these higher-order derivatives contribute only little and the CRLB becomes (approximately) reachable.
@@
In order to investigate if the effects mentioned above play a role in the mismatch between input and output FI we observe for the experimental data, we train ANNs to estimate the position of the object, each ANN with a different number of positions shown during the training. The FI is always evaluated using only the three central positions. The fraction of the input FI that arrives at the output of the ANNs trained on different parameter ranges is shown in Fig. A1.
@@
Figure A1: The fraction of the input FI that arrives at the output layer of an ANN versus the number of positions used to train the ANN. The spacing between the position is equal ($\Delta\theta=0.3$ µm); hence, more positions correspond to a broader spatial range. $13$ positions correspond to the parameter range we used to generate the results in the manuscript. We observe that the narrower the range, the more FI is conserved. The ANN layers from input to output consist of $[432,1000,500,250,50,1]$ nodes.
@@
We clearly observe that for narrower parameter ranges, more FI is conserved and when only the central three positions remain in the parameter range (those where we evaluate the FI), the mismatch disappears entirely. This result provides strong evidence that the mismatch we observed is due to the finite range of parameters we were working with.
@@
Note that there can be other reasons for such a mismatch. Examples are a (stochastic or deterministic) error in the estimated FI, a bias picked up by the estimator that cannot by taken into account with the parameter resolution at hand, or that not enough training data are available to achieve the best performance possible.
@@ A.2 Statistical properties of the estimated Fisher information
Since in our algorithm of LFI maximization we estimate the FI using a finite-sized dataset, the estimated FI itself is a random variable. Here, we provide an analytical expression for the variance of this random variable, allowing one to gauge whether the data at hand allow for a sufficiently precise estimation of the FI. Note that our estimate of the FI is the LFI of the data after it is projected into a higher dimension (see II.5). The variance of our estimate is, thus, the variance of this LFI of the transformed data. In order to approximate the variance of the LFI, we carry out a derivation similar to that in Ref. [53] and find, under the assumption that the size of the sample is large enough to accurately estimate the covariance matrix and that the sample mean, as a sum of independent random variables, approximately follows a normal distribution, the expression
@@
where $\sigma_{\hat{J}}^{2}$ is the variance of the estimated LFI $\hat{J}$, $J$ the LFI, and $d$ the dimension of the random variable and we introduce $L=JN_{\mathrm{sample}}\Delta\theta^{2}$ with the sample size $N_{\mathrm{sample}}$. To estimate the FI, we gradually increase $d$ until convergence is achieved. The results from the algorithm are precise if the fluctuations of the estimated LFI are small compared to the LFI, i.e., $\eta(L)\ll 1$ for the largest value of $d$ we chose in the algorithm. Note that if the LFI curve as a function of the dimension is not sufficiently smooth, we average over a few dimensions to smoothen the curve and identify the plateau more easily; it is reasonable to assume that this summation of LFIs does not increase the standard deviation.
From Eq. (8), we learn that the quantity $L$ should be large in order to precisely estimate the FI. This means that the sample size must be large enough to compensate for the squared finite-difference step size, which must be chosen small enough such that the discretization error remains small.
We also note that the term in Eq. (8) that is proportional to $d$ decreases with $1/L^{2}$, while the term independent of $d$ only decreases with $1/L$. Hence, for large $L$, the uncertainty of the estimated FI depends only weakly on the dimension of the projection.
@@ B.1 Optical setup
Broadband light is generated by a fiber optic illuminator (Oriel, Model No. 77501) coupled to a multimode fiber. Light is outcoupled and attenuated using a neural density filter (optical density 4) before being used to illuminate the sample. On this sample, a drawing of a space invader was engraved onto a chromium layer (CAD/Art Services). The position of the sample is controlled by a motorized stage (PI M-122.2DD1). Transmitted light is collected by a microscope objective (Olympus Plan N $\times$20, 0.4 NA) and imaged using a tube lens ($f=150$ mm) onto a camera (PCO Panda 4.2 sCMOS).
@@ B.2 Acquisition procedure
The sample is linearly translated along the x direction with a step size of $0.3$ µm and images are collected for 13 positions (measurements are taken over a broader range of positions but the outer positions are discarded to keep the data set at a moderate size). The range of positions covered by the object is, thus, comprised between x=1.8x=-1.8 µm and $x=+1.8$ µm. The image dimension is set to $32\times 64$ pixels, with a binning procedure such that each pixel corresponds to the sum of $4\times 4$ physical pixels of the camera. The exposure time is set to $10$ µs, which results in a very low signal-to-noise ratio. 300 000300\,000 images are measured for the three central positions (x=0.3x=-0.3 µm, $x=0$ µm, and $x=+0.3$ µm), and 50 00050\,000 images are measured for all other positions. Finally, a set of images is measured for different sample positions using a long exposure time ($100$ ms) for illustrative purposes.
@@ B.3 Noise of the experimental data
Figure A2: (a) shows histograms for a generic marginal distribution of the experimental data (a pixel at the center of the image). The number of histogram bins is $70$. We use a total of 50 00050\ 000 data points. (b) shows histograms for a Gaussian distribution with the same mean and variance as the experimental marginal distribution, with the same number of data points and number of bins.
@@ C.1 Generation of the Gaussian dataset for observing the Fisher information flow
80% of the dataset is used to train the ANN, while the remainder is reserved for validation. A separate dataset is prepared for evaluating the FI flow. To approximate the derivative in the expression for the FI by a symmetric finite-difference scheme, we first generate data on an equally spaced grid, but with three grid points $\theta\in\{-0.03,0,0.03\}$. 150 000 data points are generated per grid point.
@@ C.2 Generation of the experimental dataset for observing the Fisher information flow
As described in Appendix B.2, the training dataset consists of 50 00050\,000 images for each of the $13$ positions of the scatterer. Observing that information about the scatterer arrives only at a central region of the camera, we crop the images symmetrically around the center, such that the size of the images is reduced from $32\times 64$ to $18\times 24$. This procedure has a small effect on the information content of the data but reduces the memory required to store the datasets. Before it is fed to the network, the data are standardized by subtracting the intensity averaged over all images and dividing by the corresponding standard deviation. This invertible transformation does not alter the information content. 80% of the dataset is used to train the ANN while the remainder is reserved for validation.
@@
C.3 Generation of the Gaussian dataset for observing Fisher-information-based early stopping
The dataset is equivalent to that in Appendix C.1 but with $101$ parameter grid points and with $7000$ data points for each grid point and ranging from $\theta=-1.5$ to $\theta=1.5$.
C.4 Generation of the experimental dataset for observing Fisher-information-based early stopping
@@ Appendix D NEURAL NETWORK ARCHITECTURES AND TRAINING PARAMETERS
Table AI: Neural network architectures and training parameters for the ANNs employed to observe the flow of FI and early stopping based on FI, for both the Gaussian training data and the experimental training data.
@@ Appendix E CALCULATION OF FISHER INFORMATION
The LFI itself already yields satisfying results in most of the ANN’s layers. We employ the LFI maximization algorithm to find corrections to the LFI. Especially in early stages of the training there is no guarantee that the LFI is a good approximation to the true FI; thus, the LFI maximization algorithm yields more accurate and reliable results compared to using only the LFI. The linear transformations for embedding into a higher-dimensional space are random matrices with independent and identically distributed entries following a normal distribution with 0 mean and a standard deviation of $\sigma=1$. We then apply the nonlinearity
@@
with $\alpha=0.7$ to each component of the transformed random variable and eventually add noise following a normal distribution with 0 mean and a standard deviation of $\sigma_{\mathrm{noise}}$. In each iteration, we increase the dimensionality by $\Delta d$. As the criterion for convergence, we use that the difference between the smallest and the largest value of the LFI within the previous three iterations must be smaller than that of the first three iterations by a factor of $\kappa$. If during the first three iterations the increase of the LFI is smaller than a fraction of $\gamma$ of the LFI with $d^{\prime}=d$, we stop the maximization procedure and choose the LFI averaged over the first three estimates as the estimate of the true FI. The convergence criteria could not be fulfilled for all layers and epochs of the networks trained on the different datasets; thus, the LFI itself is in these cases used as a rough approximation of the true FI. If the optimized LFI is below the LFI of the dataset, we use the LFI as the estimate for the FI. This case occurs whenever the LFI is very close to the FI and the algorithm slightly underestimates the FI due to the added noise. In Table AII we list the parameters of the LFI maximization algorithm together with the exceptions, where the algorithm did not converge.
@@
Table AII: Parameters used for calculating the FI with the LFI maximization algorithm. Data are extracted at each layer of the ANNs that are trained to observe the FI flow.
@@
Note that it is not necessary to apply the LFI-based algorithm to the input layer of the ANN trained with the Gaussian training dataset as the FI of this dataset is known analytically (see Sec. II.4).
@@
Figures A3 and A4 show the LFI as a function of $d^{\prime}-d$ for an exemplary layer of the network trained on the Gaussian dataset and one of the network trained on the experimental dataset. Note that the exact shapes of these curves depend also on the magnitude of the noise that is added for regularization.
@@
Figure A3: LFI during the LFI maximization for Gaussian data. The data are extracted at the second hidden layer before the first epoch of training. The curve depicts the estimated LFI for consecutive embeddings. The dimension of the original dataset is $d=100$, and it is increased in steps of 30.
@@
Figure A4: LFI during the LFI maximization for the experimental training data set. The data are extracted at the output layer as the network is fully trained. The curve depicts the estimated LFI for consecutive embeddings. The dimension of the original dataset is $d=1$, and it is increased in steps of 30.
Appendix F PARAMETER-RESOLVED FLOW OF FISHER INFORMATION
@@
As the FI is local in the parameter-i.e., it can be different for different values of $\theta$ we can access the local flow of information through an ANN. To demonstrate this, we train an ANN on the Gaussian dataset where the estimated parameter is proportional to the mean and calculate the layerwise FI as a function of this parameter. The neural network architecture and the specifics of the training dataset are the same as those we used to observe the FI flow. We evaluate the FI at $15$ points in the range $\theta\in\{-1,1\}$ and generate for each point three samples of size 100 000100\,000, one at the point of observation $\theta$, one at $\theta-0.1$ and one at $\theta+0.1$. The result is shown in Fig. A5.
@@
Figure A5: Fisher information flow through an ANN trained on the Gaussian dataset. The curves show the (normalized) Fisher information as a function of the estimated parameter $\theta$. Each curve corresponds to a certain layer of the ANN, which is indicated by the color of the curve. The network architecture is the same as that for the flow of FI for a single parameter. Note that we stop the training at an early epoch such that it is possible to distinguish the different curves. The FI is, thus, not conserved in this example.
@@
We observe that the flow of FI is a highly nonuniform and nonsymmetric function of the parameter. The ANN, thus, learns to process data at different parameters differently. The parameters we use for our algorithm for calculating the FI are the same as in Table AII, except for $\sigma_{\mathrm{noise}}=0.05$.
@@
Appendix G TESTING OUR METHODS ON DATA FOLLOWING A CORRELATED LOG-NORMAL DISTRIBUTION